# Heart-felt Narratives: Tracing Empathy and Narrative Style in Personal Stories with LLMs

### v1

@@ Abstract
Empathy serves as a cornerstone in enabling prosocial behaviors, and can be evoked through sharing of personal experiences in stories. While empathy is influenced by narrative content, intuitively, people respond to the way a story is told as well, through narrative style. Yet the relationship between empathy and narrative style is not fully understood. In this work, we empirically examine and quantify this relationship between style and empathy using LLMs and large-scale crowdsourcing studies. We introduce a novel, theory-based taxonomy, Heart (Human Empathy and Narrative Taxonomy) that delineates elements of narrative style that can lead to empathy with the narrator of a story. We establish the performance of LLMs in extracting narrative elements from Heart, showing that prompting with our taxonomy leads to reasonable, human-level annotations beyond what prior lexicon-based methods can do. To show empirical use of our taxonomy, we collect a dataset of empathy judgments of stories via a large-scale crowdsourcing study with $N=2,624$ participants. We show that narrative elements extracted via LLMs, in particular, vividness of emotions and plot volume, can elucidate the pathways by which narrative style cultivates empathy towards personal stories. Our work suggests that such models can be used for narrative analyses that lead to human-centered social and behavioral insights.
@@
Jocelyn Shen1  Joel Mire2  Hae Won Park1 Cynthia Breazeal1Maarten Sap2,3  1Massachusetts Institute of Technology, Cambridge, MA, USA 2Carnegie Mellon University, Pittsburgh, PA, USA 3Allen Institute for AI, Seattle, WA, USA joceshen@mit.edu, jmire@andrew.cmu.edu, haewon@mit.edu, cynthiab@mit.edu, msap2@andrew.cmu.edu
@@ 1 Introduction
A key challenge in narrative analysis within the NLP community is that extracting stylistic features relevant to empathy is not trivial. Prior works use word-count-based (e.g., lexica; Roshanaei et al., 2019; Zhou et al., 2021) or hand-crafted features on extremely limited story sets Kuzmičová et al. (2017); Fernandez-Quintanilla (2020); Fernandez-Quintanilla and Stradling (2023); Eekhof et al. (2023); Mangen et al. (2018); Hartung et al. (2016) to quantify narrative elements. However, more complex stylistic narrative devices, such as plot shifts Nabi and Green (2015) or vividness of emotions Pillemer (1992) are harder to summarize with lexica alone. While a few works have explored using LLMs for more complex narrative analysis tasks Zhu et al. (2023); Michelmann et al. (2023); Sap et al. (2022), the extent to how well stylistic devices are modeled, as well as how LLM-extracted features can be used for downstream social insights has not been fully studied.
In this work, we fill this gap by presenting the following contributions. (1) We introduce Heart (Human Empathy and Narrative Taxonomy), a theory-driven taxonomy of narrative style elements that relate to empathy (Figure 2). (2) We use LLMs to quantify aspects of narrativity in our taxonomy and evaluate how well LLMs represent these elements in line with human judgments. For a subset of narrative elements with available lexica, we compare lexical measures with LLM measures, finding that in most cases, GPT-4 outperforms lexica. (3) Through a human study of $N=2,624$ participants, we introduce a new crowdsourced dataset (Heart-felt Stories Dataset) of empathetic reactions to personal narratives, including annotated narrative style elements, reader characteristics and narrative reactions. (4) With our dataset, we conduct an analysis of pathways through narrative style and reader characteristics leading to empathy, demonstrating the value of Heart in exploring empirical behavioral insights around narrative empathy. In particular, we find that narrative styles with heightened vividness of emotions, character development and action, and plot volume, are tied to narrative empathy. We additionally show that empathy is personalized, with high variability even for the same story, and that beyond narrative style, factors like a reader’s trait empathy and similarity of experiences to the narrator also significantly impact empathy.
@@ 3 Background
However, many of these narrative techniques, particularly those that are more abstract in nature, such as plot structure or emotional shifts Nabi and Green (2015), have yet to be tested empirically. Researchers in narratology have explored the impact of literary quality on reader empathy, varying aspects such as foregrounding, point of view/viewpoint words, emotion and discourse presentation, and characterisation techniques, but have found mixed results in small-scale studies Kuzmičová et al. (2017); Fernandez-Quintanilla (2020); Fernandez-Quintanilla and Stradling (2023); Eekhof et al. (2023); Mangen et al. (2018); Hartung et al. (2016). Other studies have looked at how aspects of literary reading contribute to transportation, or the ability to absorb in a narrative, which further predicts empathy towards a story Walkington et al. (2020); van Laer et al. (2014, 2019). Koopman (2015) conducted a larger-scale study to investigate the role of genre, personal factors, and affective responses on both empathic understanding and pro-social behavior, finding that genre affected prosocial behaviors. However, narrative style encompasses many aspects beyond genre alone, and each of these elements couples with one another to enhance or diminsh narrative empathy.
@@ 4 Heart Taxonomy for Empathy and Narrative Style
    Subject perception van Krieken et al. (2017) captures the vividness of perception and body sensations experienced by the character.
@@
5 Heart-felt Stories Dataset Annotation
5.1 Story Dataset
@@ 5.2 Expert Narrative Style Annotation
We randomly sampled 50 stories from our final dataset of 874 stories to obtain expert annotations of the narrative elements and validate LLM performance on the task. We selected a subset of 12 narrative elements from our taxonomy that are non-trivial to extract from existing NLP toolkits, and which required human judgments given the subjectivity of the task. Three independent members of our research team with expertise in text analysis and annotation iteratively designed a codebook (Appendix C) with instructions and examples for gauging presence of each element.
Subsequently, two independent expert annotators rated presence of each of the 12 narrative elements in the 50 sampled stories. Table 1 shows the agreement between the 2 raters using Krippendorf’s alpha (KA), percent pairwise agreement (PPA) and Spearman’s correlation ($\rho$). All ratings are positively correlated to each other, but different narrative elements have varying degrees of agreement. In particular, we observe low agreement for temporal references and perception and body sensations, where irrealis events and mentions of body sensations across multiple characters caused confusion. While some human agreements may appear low using the KA metric, these scores are consistent with prior NLP tasks with more subjectivity Shen et al. (2023); Rashkin et al. (2018); Sap et al. (2017). In our subsequent empirical analysis, we do not use features with low agreement (below $0.2$ KA).
Feature KA PPA $\rho$
Character vulnerability 62.89 86.50 $80.15^{***}$
Optimistic tone 50.97 82.25 $68.06^{***}$
Resolution 44.55 80.00 $61.59^{***}$
Character development 44.09 79.25 $61.64^{***}$
Vivid setting 42.12 78.00 $67.31^{***}$
Plot volume 33.00 79.25 $44.88^{**}$
Emotion shifts 32.25 82.25 $45.5^{**}$
Vivid emotions 27.25 75.00 $59.21^{***}$
Cognition 19.83 73.00 $34.91^{*}$
Evaluations -9.76 75.00 22.69
Table 2: Agreement between aggregated human annotators (gold ratings) and GPT-4 ratings of narrative elements in our taxonomy.
Feature $\rho_{LIWC}$ $\rho_{GPT-4}$
Optimistic tone $47.35^{***}$ $68.06^{***}$
Cognition $41.29^{**}$ $34.91^{*}$
Vivid emotions $37.63^{**}$ $59.21^{***}$
Character vulnerability -6.95 $80.15^{***}$
Table 3: GPT-4 vs LIWC narrative annotations
@@ 6 LLMs for Narrative Style Extraction
Our work explores how LLM-extracted narrative features can be used to yield empirical social insights around empathy and storytelling. As such, we validate whether LLMs are capable of narrative style annotations in line with expert human judgments. To this end, we prompt GPT-4222We used gpt-4-0613 accessed via the OpenAI API with the same instructions and codebook given to human annotators (Appendix C). In Table 2, we report agreement between averaged human ratings and GPT-4 ratings on the same 50 sampled stories.
We observe similar patterns in agreement between GPT-4 and human raters as we do in agreement between our two expert annotators. GPT-4 provides ratings with substantial agreement for narrative features such as character vulnerability, optimism, and resolution. As such, we use GPT-4 to extract the narrative elements for all the remaining stories in our corpus and exclude features that have low agreement with human gold labels in our subsequent empirical study.
@@
As prior works use lexica Roshanaei et al. (2019); Zhou et al. (2021) to quantify narrative elements, we compare whether GPT-4 can outperform psychologically validated lexica in capturing features of Heart. We select 4 dimensions in our taxonomy that readily map to lexicon-based dimensions in LIWC-22 Boyd (2022); Pennebaker et al. (1999) and compare correlation to human expert ratings in Table 3. We find that GPT-4-extracted features for optimistic tone, Vivid emotions, and Character vulnerability are better aligned with human ratings than LIWC correspondants, although only character vulnerability is statistically significantly higher ($p<0.001$ as measured by Fisher’s exact test). However, LIWC outperforms GPT-4 in the Cognition category, although not statistically significantly so. We discuss the source of potential errors in using GPT-4 to extract Cognition level of narratives in our error analyses below.
@@ 6.2 Error Analysis
Ultimately, our validation study demonstrates LLMs can approximate extracting narrative elements relevant to empathy as corroborrated by prior work Shen et al. (2023); Ziems et al. (2024), but some features are more challenging for the model to identify. We show in the following section that GPT-4 narrative ratings still reveal interesting behavioral insights around narrative empathy, even without perfect agreement.
@@ 7 Human Study for Measuring Empathy
To demonstrate empirical use of our taxonomy and how extracted narrative elements can be used to explore behavioral insights around narrative empathy, we conduct a large-scale user study presenting stories to different participants and asking them to rate their empathy towards the story. In this section, we discuss our study participants, the task procedure, and our data collection and measures used.
@@ 7.2 Study Procedure
At the beginning of the IRB-approved study, participants rated their current emotional state (arousal/valence), before reading a personal story. After reading the story, they were asked to rate their empathy towards the story, and to check which of the narrative elements within our taxonomy based on which elements contributed most to their emotional reaction towards the story. We asked a qualitative, open-ended question asking what aspects of the narrative’s style made them relate to the story.
@@ 7.3 Data Collection and Measures
Our user study aims to capture empathy towards a diverse set of narratives with a diverse set of participants with varying reader characteristics in addition to variables that might moderate the effect of narrative style on empathy. Based on related empirical work exploring factors related to empathy, we designed the following surveys (all surveys are included in Appendix E for reproducibility. We make our dataset publicly available to open up deeper research in narrative empathy analysis. [^4] URL will be made available upon publication.
@@ 8 Empirical Insights on Narrative Empathy
With annotated narrative elements, diverse empathy ratings for personal stories, and reader characteristics, we demonstrate the efficacy of our taxonomy in exploring empirical questions around empathy with a relevant subset of features from our dataset.
@@
First, we aggregate empathy ratings for each story by taking the mean across the 3 raters. Then, we split stories into high vs low presence of each narrative feature and apply Mann-Whitney u-tests to the averaged state empathy for the stories. Figure 4 shows that high aggregated empathy stories have more character development and plot volume.
@@
While our previous analysis captures aggregated empathy, different people can have diverse emotional reactions to the same story. In Figure 6 (Appendix B), we show the standard deviations in state empathy scores for the same story, finding that on average this std. dev. is significantly greater than zero ($p<0.001$), indicating that the same narrative can evoke different levels of empathy.
@@
Given our finding that narrative empathy is not “one size fits all,” we conduct analyses taking into account random effects for each story ID with structural equation modeling using the semopy library [^5] https://semopy.com/. Structural equation modeling is a standard social science method for structured hypothesis testing and uses a formulation of generalized linear models to account for fixed and random effects when a theoretical model with relationships between elements is proposed.
With our theoretical model in Figure 3, we see that vividness of emotions significantly impacts narrative transportation, which in turn influences downstream empathy towards the story. Figure 5 shows how narrative features contribute to narrative transportation, leading to downstream empathy and taking into account non-stylistic factors like the reader sharing a similar experience as the narrator and the reader’s trait empathy level. We find that both the narrator’s previous experience with something happening in the story as well as their baseline trait empathy are significant predictors of empathy towards the story, but not as much as narrative transportation.

### v2

@@ Abstract
Empathy serves as a cornerstone in enabling prosocial behaviors, and can be evoked through sharing of personal experiences in stories. While empathy is influenced by narrative content, intuitively, people respond to the way a story is told as well, through narrative style. Yet the relationship between empathy and narrative style is not fully understood. In this work, we empirically examine and quantify this relationship between style and empathy using LLMs and large-scale crowdsourcing studies. We introduce a novel, theory-based taxonomy, Heart (Human Empathy and Narrative Taxonomy) that delineates elements of narrative style that can lead to empathy with the narrator of a story. We establish the performance of LLMs in extracting narrative elements from Heart, showing that prompting with our taxonomy leads to reasonable, human-level annotations beyond what prior lexicon-based methods can do. To show empirical use of our taxonomy, we collect a dataset of empathy judgments of stories via a large-scale crowdsourcing study with $N=2,624$ participants. [^1] We make all our annotations, study data results, and language model results publicly available at https://github.com/mitmedialab/heartfelt-narratives-emnlp We show that narrative elements extracted via LLMs, in particular, vividness of emotions and plot volume, can elucidate the pathways by which narrative style cultivates empathy towards personal stories. Our work suggests that such models can be used for narrative analyses that lead to human-centered social and behavioral insights.
@@
Jocelyn Shen1  Joel Mire2  Hae Won Park1 Cynthia Breazeal1Maarten Sap2  1Massachusetts Institute of Technology, Cambridge, MA, USA 2Carnegie Mellon University, Pittsburgh, PA, USA joceshen@mit.edu, jmire@andrew.cmu.edu, haewon@mit.edu, breazeal@mit.edu, msap2@andrew.cmu.edu
@@ 1 Introduction
A key challenge in narrative analysis within the NLP community is that extracting stylistic features relevant to empathy is not trivial. Prior works use word-count-based (e.g., lexica; Roshanaei et al., 2019; Zhou et al., 2021) or hand-crafted features on extremely limited story sets Kuzmičová et al. (2017); Fernandez-Quintanilla (2020); Fernandez-Quintanilla and Stradling (2023); Eekhof et al. (2023); Mangen et al. (2018); Hartung et al. (2016) to quantify narrative elements. However, more complex stylistic narrative devices, such as plot shifts Nabi and Green (2015) or vividness of emotions Pillemer (1992) are harder to summarize with lexica alone. While a few works have explored using LLMs for more complex narrative analysis tasks Zhu et al. (2023); Michelmann et al. (2023); Sap et al. (2022), to what extent LLMs can effectively model stylistic devices, and how LLM-extracted features might be leveraged for downstream social insights, remains underexplored.
In this work, we fill this gap by presenting the following contributions. (1) We introduce Heart (Human Empathy and Narrative Taxonomy), a theory-driven taxonomy of narrative style elements that relate to empathy. (2) We use LLMs to quantify aspects of narrativity in our taxonomy and evaluate how well LLMs represent these elements in line with human judgments. For a subset of narrative elements with available lexica, we compare lexical measures with LLM measures, finding that in most cases, GPT-4 and Llama 3 outperform lexica. (3) Through a human study of $N=2,624$ participants, we introduce a new crowdsourced dataset (Heart-felt Stories Dataset) of empathetic reactions to personal narratives, including annotated narrative style elements, reader characteristics and narrative reactions. (4) With our dataset, we conduct an analysis of pathways through narrative style and reader characteristics leading to empathy, demonstrating the value of Heart in exploring empirical behavioral insights around narrative empathy. In particular, we find that narrative styles with heightened vividness of emotions, character development and action, and plot volume, are tied to narrative empathy. We additionally show that empathy is personalized, with high variability even for the same story, and that beyond narrative style, factors like a reader’s trait empathy and similarity of experiences to the narrator also significantly impact empathy.
@@ 3 Background
However, many of these narrative techniques, particularly those that are more abstract in nature, such as plot structure or emotional shifts Nabi and Green (2015), have yet to be tested empirically. Researchers in narratology have explored the impact of literary quality on reader empathy, varying aspects such as foregrounding, point of view/viewpoint words, emotion and discourse presentation, and characterisation techniques, but have found mixed results in small-scale studies Kuzmičová et al. (2017); Fernandez-Quintanilla (2020); Fernandez-Quintanilla and Stradling (2023); Eekhof et al. (2023); Mangen et al. (2018); Hartung et al. (2016). Other studies have looked at how aspects of literary reading contribute to transportation, or the ability to absorb in a narrative, which further predicts empathy towards a story Walkington et al. (2020); van Laer et al. (2014, 2019). Koopman (2015) conducted a larger-scale study to investigate the role of genre, personal factors, and affective responses on both empathic understanding and pro-social behavior, finding that genre affected prosocial behaviors. However, narrative style encompasses many aspects beyond genre alone, and each of these elements couples with one another to enhance or diminish narrative empathy.
@@ 4 Heart Taxonomy for Empathy and Narrative Style
    Subject perception van Krieken et al. (2017) captures the vividness of perception and bodily sensations experienced by the character.
@@
5 Heart-felt Stories Dataset Annotation
5.1 Story Dataset
@@ 5.2 Expert Narrative Style Annotation
We randomly sampled 50 stories from our final dataset of 874 stories to obtain expert annotations of the narrative elements and validate LLM performance on the task. We selected a subset of 12 narrative elements from our taxonomy that are non-trivial to extract from existing NLP toolkits, and which required human judgments given the subjectivity of the task. Three independent members of our research team with expertise in text analysis and annotation iteratively designed a codebook (Appendix C) with instructions and examples for gauging the presence of each element.
Subsequently, two independent expert annotators rated the presence of each of the 12 narrative elements in the 50 sampled stories. Table 1 shows the agreement between the 2 raters using Krippendorf’s alpha (KA), percent pairwise agreement (PPA), and Spearman’s correlation ($\rho$). All ratings are positively correlated to each other, but different narrative elements have varying degrees of agreement. We observe the lowest agreement between human annotators for temporal references and bodily sensations, where irrealis events and mentions of body sensations across multiple characters caused confusion. Moreover, while some human agreements may appear low using the KA metric, these scores are consistent with prior NLP tasks with more subjectivity Shen et al. (2023); Rashkin et al. (2018); Sap et al. (2017). In our subsequent empirical analysis, we do not use features with low agreement (below $0.2$ KA).
GPT-4 Llama 3 8B Instruct
Feature KA PPA $\rho$ KA PPA $\rho$
Character vulnerability 62.89 86.50 $80.15^{***}$ 27.08 79.00 70.55***
Optimistic tone 50.97 82.25 $68.06^{***}$ 48.41 82.75 67.14***
Resolution 44.55 80.00 $61.59^{***}$ 7.26 71.83 34.93*
Character development 44.09 79.25 $61.64^{***}$ 20.99 77.25 46.51**
Vivid setting 42.12 78.00 $67.31^{***}$31.07 57.03 41.84**
Plot volume 33.00 79.25 $44.88^{**}$4.00 76.08 27.51
Emotion shifts 32.25 82.25 $45.5^{**}$ 25.13 80.58 52.13***
Vivid emotions 27.25 75.00 $59.21^{***}$ 25.80 76.00 42.13**
Cognition 19.83 73.00 $34.91^{*}$ 24.98 76.00 52.89***
Evaluations -9.76 75.00 22.69 -27.16 73.00 NaN
Table 2: Agreement between aggregated human annotators (gold ratings) and GPT-4 and Llama 3 8B Instruct ratings of narrative elements in our taxonomy. Rows are sorted by GPT-4 KA.
Feature $\rho_{LIWC}$ $\rho_{GPT-4}$ $\rho_{Llama3}$
Optimistic tone $47.35^{***}$ $68.06^{***}$ $67.14^{***}$
Cognition $41.29^{**}$ $34.91^{*}$ $52.89^{***}$
Vivid emotions $37.63^{**}$ $59.21^{***}$ $42.13^{**}$
Character vulnerability -6.95 $80.15^{***}$ $70.55^{***}$
Table 3: Comparison of correlations with human annotations for LIWC, GPT-4, and Llama 3 8B Instruct.
@@ 6 LLMs for Narrative Style Extraction
Our work explores how LLM-extracted narrative features can be used to yield empirical social insights around empathy and storytelling. As such, we validate whether LLMs are capable of narrative style annotations in line with expert human judgments. To this end, we prompt GPT-4333We used gpt-4-0613 accessed via the OpenAI API. and the instruction-tuned variant of Llama 3 8B444meta-llama/Meta-Llama-3-8B-Instruct with the same instructions and codebook given to human annotators (Appendix C). In Table 2, we report agreement between averaged human ratings and the LLM-based ratings on the same 50 sampled stories.
We observe similar patterns in agreement between GPT-4 and human raters as we do in agreement between our two expert annotators. GPT-4 provides ratings with substantial agreement for narrative features such as character vulnerability, optimistic tone, and resolution. For most features, the GPT-4 ratings are more positively correlated with human annotations than are the Llama 3 ratings. As such, we use GPT-4 to extract the narrative elements for all the remaining stories in our corpus and exclude features that have low agreement with human gold labels in our subsequent empirical study.
@@
As prior works use lexica Roshanaei et al. (2019); Zhou et al. (2021) to quantify narrative elements, we compare whether GPT-4 and Llama 3 can outperform psychologically validated lexica in capturing features of Heart. We select 4 dimensions in our taxonomy that readily map to lexicon-based dimensions in LIWC-22 Boyd (2022); Pennebaker et al. (1999) and compare correlation to human expert ratings in Table 3. We find that GPT-4-extracted features for optimistic tone, vivid emotions, and character vulnerability are better aligned with human ratings than LIWC correspondents, although only character vulnerability is statistically significantly higher ($p<0.001$ as measured by Fisher’s exact test). However, LIWC outperforms GPT-4 in the Cognition category, although not statistically significantly so. We discuss the source of potential errors in using GPT-4 to extract Cognition level of narratives in our error analyses below. Notably, although Llama 3 annotations are generally relatively less correlated with human annotations, the Llama 3 extracted features consistently outperform the LIWC correspondents.
@@ 6.2 Error Analysis
Regarding Llama 3, we observe that when human annotators and GPT-4 agree, but Llama 3 disagrees, it tends to assign higher scores to a minority of features (e.g., character vulnerability) while giving lower scores to a majority of features (e.g., vivid emotions, vivid setting). The lower ratings for imagery-related features suggest a lesser adeptness with figurative language.
Ultimately, our validation study demonstrates that LLMs – in particular, GPT-4 – can approximate extracting narrative elements relevant to empathy as corroborated by prior work Shen et al. (2023); Ziems et al. (2024), but some features are more challenging for the model to identify. We show in the following section that GPT-4 narrative ratings still reveal interesting behavioral insights around narrative empathy, even without perfect agreement.
@@ 7 Human Study for Measuring Empathy
To demonstrate the empirical use of our taxonomy and how extracted narrative elements can be used to explore behavioral insights around narrative empathy, we conduct a large-scale user study presenting stories to different participants and asking them to rate their empathy towards the story. In this section, we discuss our study participants, the task procedure, and our data collection and measures used.
@@ 7.2 Study Procedure
Our study procedure was determined exempt by our institution’s ethics review board. At the beginning of the study, participants rated their current emotional state (arousal/valence), before reading a personal story. After reading the story, they were asked to rate their empathy towards the story, and to check which of the narrative elements within our taxonomy based on which elements contributed most to their emotional reaction towards the story. We asked a qualitative, open-ended question asking what aspects of the narrative’s style made them relate to the story.
@@ 7.3 Data Collection and Measures
Our user study aims to capture empathy towards a diverse set of narratives with a diverse set of participants with varying reader characteristics in addition to variables that might moderate the effect of narrative style on empathy. Based on related empirical work exploring factors related to empathy (Figure 3), we designed the following surveys (all surveys are included in Appendix E for reproducibility). We make our dataset publicly available to open up deeper research in narrative empathy analysis.
@@ 8 Empirical Insights on Narrative Empathy
Next, we demonstrate the efficacy of our taxonomy in exploring empirical questions around empathy with a relevant subset of features from our dataset.
@@
First, we aggregate empathy ratings for each story by taking the mean across the 3 raters. Then, we split stories into high vs. low presence of each narrative feature and apply Mann-Whitney u-tests to the averaged state empathy for the stories. Figure 5 shows that high aggregated empathy stories have more character development and plot volume. These results are statistically significant, after applying Benjamini-Hochberg correction to account for nine comparisons ($p=0.03$ for character development, $p=0.03$ for plot trajectory).
Our work is, to the best of our knowledge, the first to empirically test the effect of character development and plot volume on narrative empathy. While some prior works van Krieken et al. (2017) propose narrative features that relate to character identification, these are lower level than character development, such as the flatness/roundness or vulnerability of a character. Our findings regarding plot volume are in line with prior works that discuss how salient plot events can mark important moments in narratives that influence the emotional impact of the story Sap et al. (2022). Prior works primarily from narrative studies use hand-crafted features on smaller story sets Fernandez-Quintanilla (2020); Eekhof et al. (2023), but do not find significant effects of narrative features such as viewpoint and foregrounding. These studies focus primarily on literary texts rather than narratives that are more common online, and do not take into account other aspects of narrative style and narrative traits that are a part of our theorized taxonomy. These findings suggest future focused works, for example looking at how narrative style relates to empathy across narrative forms (literary vs. personal stories, spoken vs. textual, etc.)
@@
While our previous analysis captures aggregated empathy, different people can have diverse emotional reactions to the same story. In Figure 6 (Appendix B), we show the standard deviations in state empathy scores for the same story, finding that on average this std. dev. is significantly greater than zero ($p<0.001$), indicating that the same narrative can evoke different levels of empathy. To address within-subject variance, we fit mixed effects models of empathy ratings using demographic groups, grouping individuals of similar Age, Sex, Trait Empathy, and Ethnicity and conditioning on multiple ratings for a single story. We find through a likelihood ratio test that empathy predicted by demographic group results in significantly better model fit ($p=0.002$). These two results indicate that there is high variance in empathy for the same story and that incorporating information regarding diverse demographic profiles can improve empathy model fit, aligning with prior works August et al. (2020). Our findings have implications in broader empathy prediction tasks within NLP Buechel et al. (2018); Sharma et al. (2020), which often optimize for a single objective empathy score assigned to a piece of text, aggregating empathy which can overlook individual factors.
@@
Given our finding that narrative empathy is not “one size fits all,” we conduct analyses taking into account random effects for each story ID with structural equation modeling using the semopy library [^6] https://semopy.com/. Structural equation modeling (SEM) is a standard social science method for structured hypothesis testing and uses a formulation of generalized linear models to account for fixed and random effects when a theoretical model with relationships between elements is proposed.
From our SEM results (Figure 4), we find that vividness of emotions significantly impacts narrative transportation, which in turn influences downstream empathy towards the story. The importance of vividness of emotions in personal stories is supported by other work in psychology. In particular, Pillemer (1992) elaborates that vivid descriptions of emotion in personal stories can convey believability in the experience, more readily evoking empathetic responses. While some computational works explore impact of narrative features on empathy Roshanaei et al. (2019), they typically focus on positive/negative emotion words, rather than the narrative style or way in which emotions are conveyed through text, and may be better captured by current large-language models.
Figure 4 shows how narrative features contribute to narrative transportation, leading to downstream empathy and taking into account non-stylistic factors like the reader sharing a similar experience as the narrator and the reader’s trait empathy level. We find that both the narrator’s previous experience with something happening in the story as well as their baseline trait empathy are significant predictors of empathy towards the story, but not as much as narrative transportation. In particular, our findings are in line with appraisal theory that suggests that feeling similar emotions is predicated on the target sharing similar experiences (Wondra and Ellsworth, 2015; Yang and Jurgens, 2024). While it is not particularly surprising that similar experience correlates with empathy, very few works have looked at narrative style interactions in tandem with fixed (trait empathy) and more dynamic traits (experiencing something similar), suggesting more holistic consideration of contextual factors related to narrative empathy.